# Vizing’s Theorem in Near-Linear Time

### v1

@@ Appendix A Logarithmic Factors and Small Values of $\Delta$
for all $t\geq 0$. We further have,
[…]
$\leq\frac{1}{1-s\cdot(\alpha_{i}/\beta_{i})}$ (as $\beta_{i}\geq 1$)
Plugging this bounds in Equation 2, implies that
by the choice of $s$, concluding the proof. ∎
We now get to the main parts of the argument. The first part is based on a modification to the Eulerian partition approach used in the proof of Theorem 1.1.
Lemma A.6.
There is an algorithm that given a graph $G$, finds a $(\Delta+1)$ edge coloring of all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability.
Proof.
Consider the following recursive algorithm. Find an Eulerian tour of $G$ and partition $G$ into two edge-disjoint subgraphs $G_{1}$ and $G_{2}$ on the same vertex set such that $\Delta(G_{1}),\Delta(G_{2})\leq\lceil\Delta/2\rceil$ and $m(G_{1}),m(G_{2})\leq\lceil m/2\rceil$. For $i\in\{1,2\}$, recursively, find a $(\Delta(G_{i})+1)$-edge coloring $\chi_{i}$ of $G_{i}$ that leaves $m(G_{i})/\Delta(G_{i})=O(m/\Delta)$ edges uncolored. Combining $\chi_{1}$ and $\chi_{2}$ and uncoloring the two smallest color classes, gives a $(\Delta+1)$ coloring $\chi$ of all but $\lambda:=O(m/\Delta)$ edges. We then run Lemma A.4 to reduce the number of uncolored edges to $\lambda_{0}:=m/\Delta$ edges.
The correctness of the algorithm follows immediately. For the runtime, we have $k:=O(\Delta)$ sub-problems in total, with $2^{d}$ subproblems on $m/2^{d}$ edges and maximum degree $\Delta/2^{d}$ for $d\leq\left\lceil\log{\Delta}\right\rceil$. Let $X_{1},\ldots,X_{k}$ denote the runtime of these subproblems and thus $X:=\sum_{i=1}^{k}X_{i}$ is the total runtime. We have,
$\mathbb{E}\left[X\right]=\sum_{i}\mathbb{E}\left[X_{i}\right]=\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot O\left(\frac{m}{2^{d}}\log{(\lambda/\lambda_{0})}+\frac{\Delta}{2^{d}}\cdot\lambda\right)=O(m\log{\Delta})$
by Lemma A.4, since $\lambda=O(m/\Delta)$, $\lambda_{0}=m/\Delta$, and thus $\log{(\lambda/\lambda_{0})}=O(1)$ and $\Delta\cdot\lambda=O(m)$. Moreover, by Lemma A.4, for the sub-problem corresponding to $X_{i}$ at some level $d$ of the recursion, $T^{d}_{0}=O(m/2^{d})$, and every $\delta\geq 0$,
@@
For this $X_{i}$, define $\alpha_{i}:=2T^{d}_{0}$ and $\beta_{i}:=m/200\Delta$. We can now apply Proposition A.5 and for $t:=1000\sum_{i=1}^{k}\alpha_{i}=1000\cdot\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot 2T^{d}_{0}=O(m\log{\Delta})$, have,

### v3

@@ 9 Implementation and Data Structures
Consider the following recursive algorithm. Find an Eulerian tour of $G$ and partition $G$ into two edge-disjoint subgraphs $G_{1}$ and $G_{2}$ on the same vertex set such that $\Delta(G_{1}),\Delta(G_{2})\leq\lceil\Delta/2\rceil$ and $m(G_{1}),m(G_{2})\leq\lceil m/2\rceil$. For $i\in\{1,2\}$, recursively, find a $(\Delta(G_{i})+1)$-edge coloring $\chi_{i}$ of $G_{i}$ that leaves $m(G_{i})/\Delta(G_{i})=O(m/\Delta)$ edges uncolored. Combining $\chi_{1}$ and $\chi_{2}$ and uncoloring the two smallest color classes, gives a $(\Delta+1)$ coloring $\chi$ of all but $\lambda:=O(m/\Delta)$ edges. We then run Lemma 10.2 to reduce the number of uncolored edges to $\lambda_{0}:=m/\Delta$ edges. The correctness of the algorithm thus follows immediately.
For the runtime, we have $k:=O(\Delta)$ sub-problems in total, with $2^{d}$ subproblems on $\left\lceil m/2^{d}\right\rceil$ edges and maximum degree $\left\lceil\Delta/2^{d}\right\rceil$ for $d\leq\left\lceil\log{\Delta}\right\rceil$. Let $X_{1},\ldots,X_{k}$ denote the runtime of these subproblems and thus $X:=\sum_{i=1}^{k}X_{i}$ is the total runtime. We have,
$\mathbb{E}\left[X\right]=\sum_{i}\mathbb{E}\left[X_{i}\right]=\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot O\left(\frac{m}{2^{d}}\cdot\log{(\lambda/\lambda_{0})}+\frac{\Delta}{2^{d}}\cdot\lambda\right)=O(m\log{\Delta})$
by Lemma 10.2, since $\lambda=O(m/\Delta)$, $\lambda_{0}=m/\Delta$, and thus $\log{(\lambda/\lambda_{0})}=O(1)$ and $\Delta\cdot\lambda=O(m)$. Moreover, by Lemma 10.2, for the sub-problem corresponding to $X_{i}$ at some level $d$ of the recursion, $T^{d}_{0}=O(m/2^{d})$, and every $\delta\geq 0$,
@@
For this $X_{i}$, define $\alpha_{i}:=2T^{d}_{0}$ and $\beta_{i}:=m/200\Delta$. We can now apply Proposition 10.3 and for
$t:=1000\sum_{i=1}^{k}\alpha_{i}=1000\cdot\sum_{d=1}^{\left\lceil\log{\Delta}\right\rceil}2^{d}\cdot 2T^{d}_{0}=O(m\log{\Delta}),$
have,